Factor 15x^2+51X+18
I came up with (X+6) and (X-9) and then factored it to X^2-54+6X-54 but wasnt sure how to finsih it or if it was right

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Factor 15x^2+51X+18
------------
= 3(5x^2 + 17x + 6)
= 3(5x + 2)(x + 3)

Answer by stanbon(75887)

Factor 15x^2+51X+18
= (x+3)(15x+6)
=====================
Cheers,
Stan H.